### v1

# Adapting Multilingual Embedding Models to Historical Luxembourgish

Abstract
The growing volume of digitized historical texts requires effective semantic search using text embeddings. However, pre-trained multilingual models, typically evaluated on contemporary texts, face challenges with historical digitized content due to OCR noise and outdated spellings. We explore the use of multilingual embeddings for cross-lingual semantic search on historical Luxembourgish, a low-resource language. We collect historical Luxembourgish news articles spanning various time periods and use GPT-4o to segment and translate them into closely related languages, creating 20,000 parallel training sentences per language pair. We further create a historical bitext mining evaluation set and find that these models struggle to perform cross-lingual search on historical Luxembourgish. To address this, we propose a simple adaptation method using in-domain training data, achieving up to 98% accuracy in cross-lingual evaluations. We release our adapted models and historical Luxembourgish-German/French bitexts to support further research.
Adapting Multilingual Embedding Models to Historical Luxembourgish
Andrianos Michail University of Zurich andrianos.michail@cl.uzh.ch Corina Julia Raclé University of Zurich corinajulia.racle@uzh.ch
Juri Opitz University of Zurich jurialexander.opitz@cl.uzh.ch Simon Clematide University of Zurich simon.clematide@cl.uzh.ch
1 Introduction
The ability to access and explore historical texts such as newspapers has advanced rapidly through the digitization efforts of libraries and archives. This exploration has traditionally been accomplished through keyword-based searches, enhanced by semantic enrichment such as named entities.
More recently, vector-based semantic search has gained prominence through embedding models, which are also widely used in Retrieval-Augmented Generation to increase response factuality. In embedding benchmarks (Muennighoff et al., 2023), massively multilingual embedding models such as Multilingual E5 and GTE Multilingual (Zhang et al., 2024; Wang et al., 2024), trained on diverse multilingual corpora, have demonstrated strong performance in multilingual and cross-lingual semantic search across a wide range of languages. However, for low-resource languages such as Luxembourgish (LB), where multilingual models have limited exposure, their performance remains uncertain. When applied to semantic search in imperfectly digitized historical collections, these models must be robust to OCR errors and old spellings. The mismatch between this domain and the clean, contemporary texts on which these multilingual models were trained, along with their already limited support for Luxembourgish, poses a challenge for developing effective exploration tools for historical Luxembourgish newspaper archives.
To address this issue, we compile 2,340 historical Luxembourgish news articles from different time periods and use GPT-4o to segment and translate them into contemporary German (DE) and French (FR). The resulting parallel sentences serve as fine-tuning data to adapt existing multilingual embedding models for imperfectly digitized historical Luxembourgish.
Our contributions:
1. To adapt multilingual embeddings for digitized historical Luxembourgish, we propose to build training data through a prompt-based translation approach with GPT-4o.
2. For evaluation, we frame a historical Bitext Mining tasks and create a high-quality test set containing 233 source news articles (LB-DE: 2,127 sentences; LB-FR: 2,157 sentences).
3. To address this task and assess our proposed adaptation method, we evaluate models based on LaBSE (Feng et al., 2022), M-GTE (Zhang et al., 2024) and LuxEmbedder (Philippy et al., 2025), and analyze potential trade-offs in performance.
We will release our code and datasets publicly.
2 Related Work
The multilingual bitext mining model LaBSE (Feng et al., 2022) is trained with translation ranking loss and negative samples. It has been trained roughly on less than 100 Luxembourgish-English sentence pairs and specializes in zero-shot bitext mining. A recent model is GTE Multilingual (M-GTE) (Zhang et al., 2024), a multilingual embedding model designed for long context text representation and reranking. M-GTE has been trained with hard negatives and has included 50,000 Luxembourgish pairs within its contrastive pre-training.
Specific model adaptations to Luxembourgish have also been developed. One example is LuxemBERT (Lothritz et al., 2022), a monolingual BERT model pre-trained for Luxembourgish using augmented data, partially generated by translating texts from closely related languages and incorporating relevant text sources.
Closely related to our work, LuxEmbedder (Philippy et al., 2025) used OpenAI’s text-embedding-3-small and LaBSE to bitext mine a set of 20,000 parallel sentences for each language pair between Luxembourgish, English, and French. These parallel sentences were then used to further fine-tune LaBSE, improving performance on contemporary Luxembourgish evaluation sets. However, its ability to handle Luxembourgish texts from different historical periods—potentially affected by digitization errors common in large-scale historical text collections—remains unclear.
3 Method
For training and evaluating embedding models for digitized historical Luxembourgish news articles, we create parallel texts.
3.1 Parallel Historical Luxembourgish
Model Historical Bitext Mining Contemporary LB Domain Adapted LBDE LBFR SIB 200 (LB) PARALUX OpenAI/text-embedding-3-small – 82.33 78.36 40.20 15.71 OpenAI/text-embedding-3-large – 88.15 86.18 58.82 26.28 LaBSE – 94.01 93.34 43.24 38.14 LuxEmbedder – 85.48 85.07 65.59 52.24 M-GTE – 87.55 85.71 55.78 70.51 LaBSE (ours) LBDE 98.52 97.98 34.02 14.70 LBFR 98.24 97.88 35.88 25.96 LBDE/FR 98.73 98.21 36.27 20.19 LuxEmbedder (ours) LBDE 98.38 97.72 50.10 25.96 LBFR 97.89 97.74 55.88 25.32 LBDE/FR 98.52 97.91 43.04 24.68 M-GTE (ours) LBDE 97.34 96.25 44.41 41.99 LBFR 96.26 95.78 59.71 50.32 LBDE/FR 97.37 96.30 49.22 44.55 Random Baseline – 0.00 0.00 14.28 50.00
We build our translated parallel data sets LB-DE and LB-FR from monolingual Luxembourgish texts sourced from the publicly available BNL newspaper archive.111https://data.bnl.lu/data/historical-newspapers/ Our data consists of articles from newspapers published between 1841 and 1948. To select diverse samples for translation, we first cluster the articles into 2,000 groups by K-Mean and keep the 605 clusters with more than 20 articles. We select articles through a two-step process, resulting in a total of 2,340 articles, as shown in Figure 1. First, we retrieve the most representative article from each cluster, ensuring it contains between 5 and 20 sentences. In a second round, we randomly sample three additional articles per cluster under the same length conditions.
We prompt GPT-4o to segment historical Luxembourgish articles and generate sentence-level translation pairs separately for German and French (see Prompt 3). The model is instructed to preserve the original meaning and structure as closely as possible while reconstructing sentences affected by OCR errors that could hinder translation. This process yields approximately 22,500 sentence pairs for both LB-DE and LB-FR. Notably, the model’s consistent sentence segmentation results in 74.9% of sentences forming triplets across the three languages.
To ensure fidelity to the original articles, we calculate the percentage of regenerated Luxembourgish sentences that do not exactly match their source texts. These account for 1.5% of all historical Luxembourgish sentences per language pair, which we manually correct. Most mismatches result from missing or added punctuation, modernized spelling, and, in rare cases, errors caused by the LLM not adhering to the instructed format.
To assess translation quality, a trilingual native LB speaker annotates 100 randomly sampled sentence triplets after removing 15 samples with obvious OCR issues or other factors. Of 100 Luxembourgish sentences (presented without context), 86 are rated comprehensible or at least confidently guessable (24). 14 are incomprehensible in isolation due to OCR errors and old spellings. The translations of the incomprehensible sentences are not assessed. The German translations of the comprehensible and confidently guessable sentences are rated as adequate in 77 cases (89.5%): having missing details in 8 cases and only 1 case of insufficient translation. For the French translations, 76 cases are adequate (88.4%), 9 are missing minor details, and 1 is of insufficient quality. Samples from our dataset including their annotation labels are available in the Appendix (Table 2).
3.2 Framing an Evaluation Task: Historical Bitext Mining
From our parallel dataset, we reserve a held-out test set of 233 articles (2,127 sentences) to construct a historical LB-to-DE/FR Bitext Mining evaluation set. The evaluation method requires the model to assign higher similarity to the correct parallel sentence than to any of the 2,126 other candidates in the set in order to count as a true positive. We report bi-directionally averaged accuracy. To minimize false negatives stemming from near-identical sentences, we exclude candidate sentences with a Levenshtein similarity score above 0.85 to the source sentence after removing non-alphanumeric characters from both. This filtering affects 56 source-candidate pairs (2.6% of the source sentences) in German and 64 source-candidate pairs (2.97% of the sentences) in French. A human check of the excluded pairs at different thresholds confirms the appropriateness of this filtering process and the chosen threshold.
3.3 Contemporary LB Evaluation Datasets
To contextualize the evaluation on our targeted in-domain data, we replicate two evaluation methods of Philippy et al. (2025) for contemporary Luxembourgish. First, PARALUX is a monolingual Luxembourgish paraphrase detection test set designed to assess embedding models. Performance is measured by how often an embedding model assigns higher similarity to an anchor-positive pair than to an anchor-negative pair. The negative sentences are adversarially generated to ensure high lexical similarity, and human-corrected to ensure that they are true negatives. Second, SIB 200 (LB) is a repurposed dataset sourced from the ‘Flores’-dataset (Costa-jussà et al., 2022), framed as a monolingual zero-shot classification task based on embedding similarity. For further details, we refer to Philippy et al. (2025).
3.4 Adapting Multilingual Embedding Models to Historical Luxembourgish
We take the aforementioned LaBSE, LuxEmbedder and M-GTE and adapt them to historical Luxembourgish. We fine-tune these embedding models using a MultipleNegativesRankingLoss (Henderson et al., 2018) with batch size 8 over 1 epoch. We experiment with three fine-tuning data variants: (1) LBDE with 20,000 historical Luxembourgish sentences paired with German, (2) LBFR pairs, and (3) LBDE/FR, using 10,000 sentences from each language in mixed batches.
4 Results
We show the results of various models in Table 1.
Historical Bitext Mining – Cross-Lingual Semantic Search
Among off-the-shelf models, LaBSE is the strongest model in both languages. Surprisingly, LuxEmbedder, the Luxembourgish-tuned version of LaBSE, shows an 8.5pp performance drop for both language pairs in our bitext mining task, despite improved performance on contemporary Luxembourgish tasks. Similarly, M-GTE underperforms LaBSE by 5.5pp. Both OpenAI embedding models demonstrate moderate performance in all tasks.
For domain-adapted models, performance improves significantly in historical bitext mining for both language pairs after training on 20,000 parallel sentences, whether using LBDE or LBFR data. However, the best results are consistently achieved with the 20,000 LBDE/FR sentence pairs, yielding higher performance across all models. Notably, after domain adaptation, LuxEmbeddder nearly matches the adapted LaBSE, achieving over 98% accuracy and closing the performance gap observed in the off-the-shelf models. In contrast, M-GTE, the general embedding model, appears to plateau at lower performance levels, trailing the specialized domain-adapted bitext mining model by 1.5pp.
Contemporary vs. historic Luxembourgish: A notable trade-off
While our domain-adapted models improve the embedding of historical Luxembourgish for cross-lingual semantic search in contemporary German and French (as shown on the right side of the Table 1), their performance on monolingual contemporary tasks, particularly PARALUX, declines noticeably after domain adaptation. Interestingly, this degradation is more pronounced when models are trained on the linguistically closer LBDE pairs than LBFR pairs.
The learning curves in Figure 2 illustrate the continuous decline in monolingual performance. Notably, most cross-lingual performance gains occur within the first 2,500 samples, before monolingual Luxembourgish performance declines more sharply. This highlights the trade-off between optimizing for historical cross-lingual tasks and maintaining monolingual performance, as seen in the weaker monolingual results of the best off-the-shelf LaBSE model.
5 Conclusions
In this work, we present a case study on embedding imperfectly digitized historical Luxembourgish texts. As off-the-shelf models prove suboptimal for this task, we propose a method to synthesize reliable data for adapting embedding models to this domain. We generate parallel sentence-segmented documents by translating historical Luxembourgish newspaper articles into two related higher-resource languages (German and French) using GPT-4o as a machine translation system.
To establish a hold-out test set, we reserve 233 articles and frame a bitext mining task for retrieving historical Luxembourgish sentences. Evaluating various multilingual embedding models on this task, we demonstrate that our adaptation approach improves performance by up to 13 percentage points. While our adapted models perform effectively across different historical periods, this improvement comes at the cost of reduced performance in natural language understanding (NLU) for clean contemporary texts. This trade-off highlights the need for future research into embedding historical digitized texts affected by significant OCR errors.
Limitations
Our findings present both positive and negative aspects. On the one hand, our method demonstrates clear benefits for the targeted use case, effectively embedding heterogeneous digitized historical texts and revealing shortcomings in off-the-shelf models. On the other hand, performance drops in control tasks on contemporary texts raise concerns. Does improved robustness to digitization errors come at the cost of weaker language understanding and reduced sensitivity to semantic nuances? Can future training strategies mitigate this trade-off? These important questions remain open for future research.
Acknowledgements
This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
Appendix A Appendix
| Newspaper | Year | Sentence | LB Comprehensibility Rating | DE Translation Rating | FR Translation Rating |
| D’Union | 1946 |
LB: Si bestét aus 18 000 tuben, weit 30 tonnen a kascht 400 000 dollar.
DE: Sie besteht aus 18.000 Röhren, wiegt 30 Tonnen und kostet 400.000 Dollar. FR: Elle est composée de 18 000 tubes, pèse 30 tonnes et coûte 400 000 dollars. |
Comprehensible | Adequate | Adequate |
| Ons Jorgen | 1946 |
LB: „Daf net, mais ech hu kc properen Teller me’,"
DE: „Das nicht, aber ich habe keinen sauberen Teller mehr. FR: "Certainement pas, mais je n’ai plus d’assiette propre," |
Comprehensible | Adequate | Adequate |
| De Letzeburger | 1893 |
LB: De Batti: Elo hätte mer d’Stemmung gut eriwer, hätte mer elo och nach Rén.
DE: Der Batti: Jetzt hätten wir die Stimmung gut geschafft, hätten wir jetzt auch noch Regen. FR: Le Batti : Maintenant, nous aurions bien passé l’ambiance, si seulement nous avions aussi encore Rén. |
Confidently Guessable | Adequate | Adequate |
| De Letzeburger | 1905 |
LB: Op d’Weis: Das ist im Lehen hfisslich eingerichtet.
DE: Zur Melodie: Das ist im Leben hässlich eingerichtet. FR: À la manière de : Cela est arrangé vilaine dans la vie. |
Confidently Guessable | Adequate | Adequate |
| De Letzeburger | 1893 |
LB: Wann och d’Liss’ché wéss ze feischtren.
DE: Wenn auch die Liss’ché weiß zu feilschen. FR: Même si Liss’ché sait lutter. |
Incomprehensible | / | / |
| Obermosel-Zeitung | 1924 |
LB: In vielen vorkern Bincl alle Krank, 80 cla,BB «lie ?eläer nicdt deBtellt terrien Können.
DE: In vielen Dörfern sind alle krank, so dass die Felder nicht bestellt werden können. FR: Dans de nombreux villages, tous sont malades, si bien que les champs ne peuvent pas être cultivés. |
Incomprehensible | / | / |